order: intropin 0.5mcg/kg/min for renal vasodilation

instructions: add 100mg to 1000 mL

Drop factor : 15gtts/mL

Clients weight: 212 lbs

Calculate the flow rate?
Answer #1

ANSWER.

* Calculation of flow rate.

      Formula

         Flow rate(gtt/min) = Volume(ml) /Time(min) * Drop factor(gtts/ml).

    Calculation of volume(ml).

         Formula = Desired dose/Dose in hand * Quantity.

Here

         Ordered dose = 0.5 mcg/kg/min.

       Convert mcg to mg for calculation

                  1 mcg = 0.001 mg

      So

          Ordered dose = 0.5 * 0.001

                                  = 0.0005 mg/kg/min.

         Weight of the patient = 212 lbs.

     Convert lbs to kg for calculation

                1 lbs = 0.45 4 kg.

       Weight of the patient = 212 * 0.454

                                         = 96.2 kg.

         Desired dose for this patient(D) = 0.0005 * 96.2

                                                            = 0.05 mg/min.

                 Dose in hand (H) = 100 mg.

                    Quantity(Q) = 1000 ml.

   As per above formula for calculation of volume(ml).

               Volume(ml) = D/H *Q

                                   = 0.05 / 100 * 1000

                                   = 0.0005 * 1000

                                   = 0.5 ml.

          Flow rate = Volume/Time* Drop factor.

               Here

                   Volume(ml) = 0.5 ml.

                     Time (min) =1 min.

                   Drop factor = 15 gtt/ml.

        So

             Flow rate(gtt/min) = 0.5/1 * 15

                                            = 7.5 gtt/min.